OpenStudy (anonymous):

Find the distance between the points (-4, 6) and (8, -12).Round your solution to 2 decimal points. can someone teach me how to solve this kind of problem. im very confused.

OpenStudy (anonymous):

d=sqrt[{8-(-4)}^2+(-12-6)^2] =sqrt(144+324) =sqrt468 =21.63

OpenStudy (anonymous):

thanks :) so just so i have this in formula, is D=(x2-x1)^2+(y2-y1)^2 correct

OpenStudy (anonymous):

D=sqrt(x2-x1)^2+(y2-y1)^2 ops forgot the sqrt

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thanks for the help :)